A wheel has a diameter 28 inches. If it turns ten times, how far has the vehicle traveled, to the nearest foot?: A wheel has a diameter 28 inches. If it turns ten times, how far has the vehicle traveled, to the nearest foot?
Answer by orca(336)

Formula to use:
The perimeter of a circle with diameter D is pi*D
When the wheel turns one time, the vehicle will move a distance of pi*D, If it turns 10 times, the vehicle will cover a distance of 10*pi*D.
Therefore, The distance the vehicle traveled is:
10*pi*D
=10*3.14*28
= 879.2 inches
Converting it to foot,we have(Note that 1 feet = 12 inches, so 1 inch = 1/12 feet)
879.2 inches = 879.2*1 inches = 879.2*(1/12)foot=879.2/12 foot = 73.3 foot